OpenStudy (anonymous):

int_{?}^{?} (1/2y - 2/y^2 + 3/y^1/2) dy

OpenStudy (anonymous):

\[\int\limits_{?}^{?} (1\div2y - 2/y^2 + 3/\sqrt{y}) dy\]

OpenStudy (anonymous):

CThis integral looks like that to me \[\int\limits_{}^{}(1/(2y)) -(2/y^3)+(3/\sqrt{y})dy\]. If that is the case then you can separate the integral into three integrals, and bring the y's to the numerator by making the exponent negative. This makes it easier to view when resolving. the only one that will not do that is the first one, which will boil down to (1/2) ln y after integrating. bring your constants in front of the integrals and bring the exponents to their appropriate negative power. Remember the first one is a 1/y situation = ln